# The AIDA-TNG project: 3D halo shapes

### v1

@@ Abstract
Context. The shapes of dark matter halos can be used to constrain the fundamental properties of dark matter. In standard Cold Dark Matter (CDM) cosmologies, halos are typically triaxial, with a preference for prolate configurations, particularly at low masses and high redshift.
@@
Methods. We measure the different halo component shapes – dark matter, stars and gas – at various radii from the center in the AIDA-TNG (Alternative Interacting Dark Matter and Astrophysics – TNG), which is a suite of high-resolution cosmological simulations built upon the IllustrisTNG framework. The intent is to systematically study how different dark matter models – specifically, SIDM and WDM – affect galaxy formation and the structure of dark matter halos, when realistic baryonic physics is also included.
Results. SIDM models tend to produce rounder and more isotropic halos, especially in the inner regions, as a result of momentum exchange between dark matter particles. WDM halos are also slightly more spherical than their CDM counterparts, and are typically less concentrated. In all cases, the inclusion of self-consistent baryonic physics makes the central regions of all halos rounder, while still revealing clear distinctions among the various dark matter models.
@@ 1 Introduction
Substructures and anisotropic accretion further influence halo morphology (Nipoti et al. 2018). Massive satellites often preserve their infall directions along filaments, while smaller satellites are accreted from a broader range of angles, affecting halo triaxiality in the outskirts (Knebe et al. 2004; Libeskind et al. 2013a, b; Forero-Romero et al. 2014). The combined effects of baryons, substructure, and filamentary accretion produce complex, radially dependent halo shapes. Observational studies infer triaxial halo shapes through gravitational lensing, X-ray measurements, the Sunyaev-Zel’dovich effect, and the spatial distribution of satellites (De Filippis et al. 2005; Oguri et al. 2010; Sereno et al. 2013; Gonzalez et al. 2021). Typically, the observable is a projected two-dimensional shape; we defer a detailed study of the inversion of this projection into 2D detectable quantities to a separate paper.
In this work, we use the AIDA-TNG simulations (Despali et al. 2025), a suite of high-resolution cosmological magnetohydrodynamical runs, to explore halos’ three-dimensional shapes and internal properties across CDM, SIDM, and WDM scenarios. The AIDA-TNG simulations incorporate full baryonic physics, including gas cooling, star formation, and feedback processes, allowing us to quantify the interplay between baryons and DM microphysics (Vogelsberger et al. 2014a, b; Rodriguez-Gomez et al. 2016, 2017; Pillepich et al. 2018a). The reference DM-only runs will also allow us to quantify the effects of baryonic physics implementations and possible impact on DM particle properties. We focus on mass-selected halo samples across different DM models, without attempting direct halo matching, to provide a statistically representative assessment of how halo morphology responds to baryonic processes and alternative DM properties.
@@ 2 AIDA-TNG Cosmological Hydrodynamical Simulations
The AIDA-TNG simulations, Alternative Dark Matter in the TNG universe (AIDA for short, hereafter) (Despali et al. 2025), consist of three cosmological volumes of increasing resolution and decreasing size simulated in cold and alternative dark matter (ADM, hereafter) models, with and without baryonic physics. This work utilizes the largest boxes, measuring 75 $h^{-1}$ Mpc (110.7 Mpc) on a side (see Table 1 for more details). The uniqueness of the AIDA set lies in the fact that we can utilize the same cosmological volumes across multiple dark matter scenarios, encompassing cold, warm, and self-interacting dark matter. In particular, we make use of the WDM3 run, corresponding to a particle mass of 3 keV, and two different self-interacting models, namely SIDM1 with a constant cross section equal to 1 cm2 g-1, and vSIDM with a scale-dependent cross section as described in Correa (2021). Moreover, we can rely on both a dark matter-only (DMO) and a full-physics (FP) version of each run so that we can disentangle the effects, relative to a standard CDM run, of ADM and baryonic physics as described by the IllustrisTNG galaxy formation model (TNG hereafter Weinberger et al. 2017; Pillepich et al. 2018a).
@@ 3.1 Shape parameters
The weighting in the inertial tensor ensures convergence toward a self-consistent ellipsoidal shape (Jing & Suto 2002; Allgood et al. 2006; Zemp et al. 2011). Shape determination is performed iteratively: starting from a spherical volume, the tensor is computed, diagonalized, and the resulting axis ratios are used to redefine the ellipsoidal boundary, enclosing $N$ particles — also updated at each step – until a $1\%$ convergence on the axis ratios is achieved. The resulting axis ratios describe the flattening of the halo. If $a\geq b\geq c$ then, following Bardeen et al. (1986) and Matarrese et al. (1991), we define the ellipticity,
@@
We have prolate those halos where $\epsilon<p<0$; oblate halos where $0<p<\epsilon$. In the limiting cases where $p=-\epsilon$ or $p=\epsilon$, one obtains either a prolate or an oblate spheroid, respectively. We can also quantify the degree of triaxiality using the parameter
@@
Figure 2: Profiles of minor-to-major (top) and intermediate-to-major (bottom) axis ratios for $z=0$ halos in the CDM simulation in bins of increasing mass (from left to right). Solid black curves show the total axis ratios in the FP run, while solid red, dotted gold, and solid orange curves show the corresponding results for the dark matter, gas, and stars, respectively. Dashed red curves show halos in the same mass bins but for the dark matter-only run. Green data points (shown only in some of the boxes) consider the satellites and were weighted by stellar mass when computing the inertial mass tensor. The error bars bracket 25% and 75% of the distribution. The vertical bars display the resolution limit, which is equal to 10 times the softening length.
@@ 3.2 Shapes in the CDM model
In Fig. 2, we present the radial profiles of the intermediate-to-major ($b/a$, top panels) and minor-to-major ($c/a$, bottom panels) axis ratios for all mass components of halos at $z=0$ in the CDM scenario. We consider four distinct mass bins, as indicated at the top of each panel. These profiles illustrate how the three-dimensional morphology of halos varies from the virial radius down to the innermost regions, and how this behavior depends on the baryonic content and feedback processes. In the DMO run (red dashed lines), we recover the expected dependence of the axis ratios on distance, with a higher triaxiality towards the halo center (e.g. Allgood et al. 2006; Despali et al. 2017; Chua et al. 2019).
@@
Star particles display more triaxial shapes than the dark matter at nearly all radii. This enhanced triaxiality likely originates from the complex assembly history of the stellar component, which includes contributions from in-situ star formation as well as the accretion of satellite galaxies. The alignment of the stellar component with the underlying dark matter potential is therefore only partial, consistent with previous findings from hydrodynamical simulations (e.g. Tenneti et al. 2014; Chua et al. 2019). In addition is worth mentioning that baryons make haloes rounder and more oblate, and strengthen the alignment between the stellar and dark matter components (Chua et al. 2022). These effects become more pronounced with increasing stellar-to-halo mass ratio, indicating that systems with higher baryon conversion efficiency experience a stronger modification of their inner potential. This trend agrees with our AIDA–TNG results, where the inner halo becomes progressively more spherical as the baryonic mass fractions increase, highlighting the central role of feedback-regulated galaxy formation in shaping the dark matter distribution. We discuss this in more detail in the next section.
@@ 3.3 The effect of SIDM and WDM on shapes
Elastic scattering between dark matter particles transfers momentum and heat within the halo, leading to the formation of a central core and a substantial redistribution of momentum through particle scattering (e.g. Rocha et al. 2013; Spergel & Steinhardt 2000; Robertson et al. 2021). Consequently, SIDM halos are expected to be more spherical than their CDM counterparts, particularly within the inner $\sim 0.1\,R_{\mathrm{vir}}$.
SIDM halos, both for constant and scale-dependent cross sections, are more spherical than CDM in the DMO run (bottom panels of Fig. 3). This trend, albeit diminished, persists when baryons are included, although the combined effect depends on the efficiency of feedback and the characteristic cross-section of the interaction. In the top panels, we see that the baryonic and feedback mechanisms influence and reduce the differences between CDM, SIDM, and vSIDM, especially in the central regions. In all panels, the WDM3 is relatively close, within 5%, to the CDM case. The small differences are due to the small-scale suppression, late forming, and less concentrated WDM halos.
@@ 4 Misalignment angles between star, gas, and dark matter components
Figure 4: Median misalignment angles of dark matter (red) and gas (orange) with respect to the star particles (yellow), in the AIDA-TNG simulations. Dashed, dot-dashed, and dotted lines refer to SIDM1, vSIDM, and WDM3 simulations, respectively, and the four panels are for the same mass bins as in Fig. 2. The colored shaded regions bracket the central 50% (i.e., between 25% and 75%) of the CDM measurements. The four panels refer to various mass bins as presented in Fig. 2
In the AIDA-TNG simulations, we compute the misalignment between components by measuring the angle between the major axes of the corresponding inertia tensors, derived for the stellar with respect to the dark matter and gas particle distributions. This is done in light of considering optically selected stellar components from photometric observations. This approach allows us to quantify the three-dimensional orientation of each component as a function of radius and halo mass. Results are displayed in Fig. 4. Red, orange, and gold lines represent the dark matter, gas, and stellar components of various halos and dark matter models. CDM, SIDM1, vSIDM, and WDM3 are represented by solid, dashed, dot-dashed, and dotted lines, respectively. The four panels refer to various mass bins as presented in Fig. 2.

### v2

@@ Abstract
Context. The shapes of dark matter halos can be used to constrain the fundamental properties of dark matter. In standard Cold Dark Matter (CDM) cosmologies, halos are typically triaxial, with a preference for prolate configurations; however, including the full baryonic physics tends to make them more oblate.
@@
Methods. We measure the different halo component shapes – dark matter, stars, and gas – at various radii from the center in the AIDA-TNG (Alternative Interacting Dark Matter and Astrophysics – TNG), which is a suite of high-resolution cosmological simulations built upon the IllustrisTNG framework. The intent is to systematically study how different dark matter models – specifically, SIDM and WDM – affect galaxy formation and the structure of dark matter halos, when realistic baryonic physics is also included.
Results. SIDM models tend to produce rounder and more isotropic halos, especially in the inner regions, as a result of momentum exchange between dark matter particles. Group and cluster-size WDM halos are also slightly more spherical than their CDM counterparts. In all cases, the inclusion of self-consistent baryonic physics makes the central regions of all halos rounder, while still revealing clear distinctions among the various dark matter models, mainly valid for the self-interacting ones.
@@ 1 Introduction
Substructures and anisotropic accretion further influence halo morphology (Nipoti et al. 2018). Massive satellites often preserve their infall directions along filaments, while smaller satellites are accreted from a broader range of angles, affecting halo triaxiality in the outskirts (Knebe et al. 2004; Libeskind et al. 2013a, b; Forero-Romero et al. 2014). The combined effects of baryons, substructure, and filamentary accretion produce complex, radially dependent halo shapes. Observational studies infer triaxial halo shapes through gravitational lensing, X-ray measurements, the Sunyaev-Zel’dovich effect, and the spatial distribution of satellites (De Filippis et al. 2005; Oguri et al. 2010; Sereno et al. 2013; Gonzalez et al. 2021). Typically, the observable is a projected two-dimensional shape (Meneghetti et al. 2007b); we defer a detailed study of the inversion of this projection into 2D detectable quantities to a separate paper.
In this work, we use the AIDA-TNG simulations (Despali et al. 2025b), a suite of high-resolution cosmological magnetohydrodynamical runs, to explore halos’ three-dimensional shapes and internal properties across CDM, SIDM, and WDM scenarios. The AIDA-TNG simulations incorporate full baryonic physics, including gas cooling, star formation, and feedback processes, allowing us to quantify the interplay between baryons and DM microphysics (Vogelsberger et al. 2014a, b; Rodriguez-Gomez et al. 2016, 2017; Pillepich et al. 2018a; Vogelsberger et al. 2020). The reference DM-only runs will also allow us to quantify the effects of baryonic physics implementations and possible impact on DM particle properties. We focus on mass-selected halo samples across different DM models, without attempting direct halo matching, to provide a statistically representative assessment of how halo morphology responds to baryonic processes and alternative DM properties.
@@ 2 AIDA-TNG cosmological hydrodynamical simulations
The AIDA-TNG simulations, Alternative Dark Matter in the TNG universe (AIDA for short, hereafter) (Despali et al. 2025a, b; Romanello et al. 2025), consist of three cosmological volumes of increasing resolution and decreasing size simulated in cold and alternative dark matter (ADM, hereafter) models, with and without baryonic physics. This work utilizes the largest boxes, measuring 75 $h^{-1}$ Mpc (110.7 Mpc) on a side (see Table 1 for more details). The uniqueness of the AIDA set lies in the fact that we can utilize the same cosmological volumes across multiple dark matter scenarios, encompassing cold, warm, and self-interacting dark matter. In particular, we make use of the WDM3 run, corresponding to a particle mass of 3 keV, and two different self-interacting models, namely SIDM1 with a constant cross section equal to 1 cm2 g-1, and vSIDM with a scale-dependent cross section as described in Correa (2021). Moreover, we can rely on both a dark matter-only (DMO) and a full-physics (FP) version of each run so that we can disentangle the effects, relative to a standard CDM run, of ADM and baryonic physics as described by the IllustrisTNG galaxy formation model (TNG hereafter Weinberger et al. 2017; Pillepich et al. 2018a).
@@ 3.1 Shape parameters
The weighting in the inertial tensor ensures convergence toward a self-consistent ellipsoidal shape (Jing & Suto 2002; Allgood et al. 2006; Zemp et al. 2011). Shape determination is performed iteratively: starting from a spherical volume, the tensor is computed, diagonalized, and the resulting axis ratios are used to redefine the ellipsoidal boundary, enclosing $N$ particles — also updated at each step – until a $1\%$ convergence on the axis ratios is achieved. The resulting axis ratios describe the flattening of the halo (Bardeen et al. 1986; Matarrese et al. 1991). If $a\geq b\geq c$, we define the ellipticity,
@@
We have prolate those halos where $p>0$; oblate halos where $p<0$. In the limiting cases where $p=\epsilon$ or $p=-\epsilon$, one obtains either a prolate or an oblate spheroid, respectively. We can also quantify the degree of triaxiality using the parameter
@@
Figure 2: Median profiles of minor-to-major (top) and intermediate-to-major (bottom) axis ratios for $z=0$ halos in the CDM simulation in bins of increasing mass (from left to right). Solid black curves show the total axis ratios in the FP run, while solid red, dotted gold, and solid orange curves show the corresponding results for the dark matter, gas, and stars, respectively. Dashed red curves show halos in the same mass bins but for the DMO run. Green data points (shown only in some of the boxes) consider the satellites and were weighted by stellar mass when computing the inertial mass tensor. The error bars bracket 25% and 75% of the distribution. The vertical bars display the resolution limit, which is equal to 10 times the softening length. The number of halos present in the four considered mass bins is: 305, 78, 34, and 14, respectively.
@@ 3.2 Shapes in the CDM model
In Fig. 2, we present the median radial profiles of the intermediate-to-major ($b/a$, top panels) and minor-to-major ($c/a$, bottom panels) axis ratios for all mass components of halos at $z=0$ in the CDM scenario. We consider four distinct mass bins, as indicated at the top of each panel, containing 305, 78, 34, and 14 systems ordered in increasing mass. These profiles illustrate how the three-dimensional morphology of halos varies from the virial radius down to the innermost regions, and how this behavior depends on the baryonic content and feedback processes. In the DMO run, displayed using a red dashed style line, we recover the expected dependence of the axis ratios on distance, with a higher triaxiality towards the halo center (e.g. Allgood et al. 2006; Despali et al. 2017; Chua et al. 2019). We underline that we use the solid red line style for the DM component of the FP run.
@@
Star particles display more triaxial shapes than the dark matter at nearly all radii. This enhanced triaxiality likely originates from the complex assembly history of the stellar component, which includes contributions from in-situ star formation as well as the accretion of satellite galaxies. The alignment of the stellar component with the underlying dark matter potential is therefore only partial, consistent with previous findings from hydrodynamical simulations (e.g. Tenneti et al. 2014; Chua et al. 2019). In addition is worth mentioning that baryons make halos rounder and more oblate, and strengthen the alignment between the stellar and dark matter components (Chua et al. 2022). These effects become more pronounced with increasing stellar-to-halo mass ratio, indicating that systems with higher baryon conversion efficiency experience a stronger modification of their inner potential. This trend agrees with our AIDA–TNG results, where the inner halo becomes progressively more spherical as the baryonic mass fractions increase, highlighting the central role of feedback-regulated galaxy formation in shaping the dark matter distribution. We discuss this in more detail in the next section.
@@ 3.3 The effect of SIDM and WDM on shapes
Elastic scattering between dark matter particles transfers momentum and heat within the halo, leading to the formation of a central core and a substantial redistribution of momentum through particle scattering (e.g. Rocha et al. 2013; Spergel & Steinhardt 2000; Robertson et al. 2021). Consequently, SIDM halos are expected to be more spherical than their CDM counterparts, particularly within the inner $\sim 0.1\,R_{\mathrm{200}}$.
SIDM halos, both for constant and scale-dependent cross sections, are more spherical than CDM in the DMO run (bottom panels of Fig. 3). This trend, albeit diminished, persists when baryons are included, although the combined effect depends on the efficiency of feedback and the characteristic cross-section of the interaction. In the top panels, we see that the baryonic and feedback mechanisms influence and reduce the differences between CDM, SIDM, and vSIDM, especially in the central regions. In all panels, the WDM3 is relatively close, within 5%, to the CDM case. The small differences are due to the small-scale suppression, late forming, and less concentrated WDM halos. For the two largest mass bins – the last two column subpanels – we can notice that at 10% the halo boundary radius, there is a clear difference between SIDM1 and vSIDM models, where the halos in the former are more spherical than in the latter. However, at smaller radii, the behavior changes. While vSIDM halos remain more spherical than their CDM counterparts, SIDM1 halos exhibit shapes that are comparable to those in CDM. These trends are qualitatively similar in both the DMO and FP runs, underlining that in the latter, the shape toward the center is more strongly influenced by baryonic physics mechanisms (Despali et al. 2022).
@@ 4 Misalignment angles between star, gas, and dark matter components
Figure 4: Median misalignment angles, as a function of the rescaled radius, of dark matter (red) and gas (orange) with respect to the star particles (yellow), in the AIDA-TNG simulations. Dashed, dot-dashed, and dotted lines refer to SIDM1, vSIDM, and WDM3 simulations, respectively, and the four panels are for the same mass bins as in Fig. 2. The colored shaded regions bracket the central 50% (i.e., between 25% and 75%) of the CDM measurements. The four panels refer to various mass bins as presented in Fig. 2
In the AIDA-TNG simulations, we compute the misalignment between components by measuring the angle between the major axes of the corresponding inertia tensors, derived for the stellar with respect to the dark matter and gas particle distributions. This is done in light of considering optically selected stellar components from photometric observations. This approach allows us to quantify the three-dimensional orientation of each component as a function of radius and halo mass. Results are displayed in Fig. 4. Red and orange lines represent the misalignment of dark matter and gas components with respect to the stars. CDM, SIDM1, vSIDM, and WDM3 are represented by solid, dashed, dot-dashed, and dotted lines, respectively. The four panels refer to various mass bins as presented in Fig. 2.